If vector A +B +C=0 and A=B=C THEN FIND ANGLE BETWEEN VECTOR A AND VECTOR B.

Grade:11

1 Answers

Eshan
5 years ago
$$\vec{A}}+\vec{B}}+\vec{C}}=0$$

$$\implies \vec{A}+\vec{B}=-\vec{C}$$
$$\implies \left|\vec{A}+\vec{B}\right|=\left| -\vec{C}}\right|$$
$$\implies A^{2}+B^{2}+2AB cos{\theta}=C^{2}$$
Since A=B=C,

We have,$$\cos{\theta}=-\dfrac{1}{2}$$
$$\implies \theta=120^{\circ}$$